# A Comparative Study of Learning Paradigms in Large Language Models via Intrinsic Dimension

### v1

@@ 7.1 Experiment Setup for Comparative Analysis
We perform a series of experiments directly comparing ID curves from both SFT and ICL. We utilize similar setups as discussed in Sections 5 and 6. However, for the fine-tuning experiments in this section, we only train for 4 epochs and only measure the accuracy and ID for the final checkpoint. We choose this number of epochs because the previous experiments in Section 5 indicated that models tended to begin over-fitting past 4 epochs across the datasets tested.
For ICL experiments in this section, we only consider 0,1,2,5,10 as values for $k$. We choose these values for the number of demonstrations because they are popular choices for ICL in practice and also because prior experiments in 6 indicate that ID curves tended to plateau when $k\geq 10$. We perform experiments on all 8 datasets and 4 models discussed in Section 4.
@@ 7.2 In-context Learning Induces Higher IDs Compared to Fine-Tuning
We find that across all datasets and models, ICL with $k\geq 5$ consistently induces higher intrinsic dimensions (IDs) across all layers compared to SFT and 0-shot prompts (see Figures 1b and 1c). This contrasts with the findings of Doimo et al. (2024), who report that ID induces higher ID in later layers.
@@
Figure 5: Boxplot displaying the distribution of normalized AUC values for different learning paradigms. Each point corresponds to the normalized AUC value for a (model, dataset) pair. Median normalized AUC peaks with 5-shot ICL, while values for SFT are closer to zero-shot baseline (icl-0).
We also find that the ID values of models fine-tuned with 1000 samples tend to remain similar to the original ID of the baseline model on a zero-shot prompt (designated by icl-0). We present a heatmap displaying average differences in normalized AUC between learning paradigms in Figure 4, and a boxplot depciting the distribution of normalized AUC values for the different paradigms in Figure 5.
@@ 7.3.1 Differing Shapes of Intrinsic Dimension Curves
We observe that the exact shape of the Intrinsic Dimension curves is highly dependent upon the dataset. For some datasets, such as AG News, we observe a concsistent ’hunchback’ shape, where ID increases is progressively lowered throughout the later layers of the model, across all models and learning paradigms (Figure 36). This shape has been reported on by previous work by Yin et al. (2024) in QA datasets. However, we find that this pattern does not consistently hold across all models, datasets, and learning paradigms. For example, on the QQP dataset, we do not observe a consistent hunchback shape for icl-0, icl-1 or fine-tune learning paradigms (33). Thus, in contrast to Convolutional Neural Networks (Ansuini et al., 2019), Image Generation Transformers such as ImageGPT and Protein Language Models (Valeriani et al., 2024), which have been found to exhibit consistent patterns Intrinsic Dimension patterns across their layers for inputs of their respective data modality. This difference suggests that LLMs encode data into more diverse manifolds in their representation space, potentially reflecting their generality and the complexity of their learning task compared to other neural networks.
Figure 6: Boxplot displaying the distribution of normalized AUC values across datasets for each model in the ICL-5 shot setting. Each point corresponds to the value corresponding to a (model, dataset) pair. ID values all lie in a narrow range, highlighting similarity in representation spaces across models.
We also find that for a specific learning paradigm, the range of normalized AUC values across datasets for the four different models we tested to be in similar ranges, despite the fact that the models come from different families and have different embedding dimensions (e.g. Llama-2-13b has a hidden dimension of 5120, while the other three models have hidden dimensions of 4096). Figure 6 depicts the range of normalized AUC values for the icl-5 learning paradigm and shows that all the values fall within a range of 20. We view this as evidence that different models may be generating representations with similar geometric complexity for a specific dataset, despite differences in model size or pre-training schemes. We include similar boxplots for normalized AUC values of other experiments in Appendix E. These findings are in agreement with results from Cheng et al. (2023), which find that LLMs of different sizes and families create representations with similar ID values for a variety of text corpora.
@@ 7.4 Comparing Performance of Different Learning Paradigms
We found that models fine-tuned with 1k samples obtained the highest accuracy, and models performing ICL with 10-samples followed closely. This implies ID may not be closely related to accuracy, as the models fine-tuned with 1k samples had a closer ID to the base models than to ICL models, which had substantially lower accuracies. See Table 1 for average performance of each of the learning paradigms across the models and datasets tested.
@@ 8 Summary
We present a detailed analysis of the intrinsic dimension (ID) induced by the SFT and ICL learning paradigms. We identify a trend where the normalized AUC of ID curves sometimes decreases during the initial stages of SFT but generally increases during the later stages of fine-tuning.
Additionally, we observe a trend in the number of demonstrations used in ICL and the resulting ID curves: the normalized AUC of the ID curves initially increases for small values of $k$ (the number of demonstrations) but plateaus or slightly decreases as $k$ increases further. We observe that the value of $k$ which corresponds to the highest normalized AUC also obtains peak (or close-to-peak) accuracy, suggesting that ID may serve as a useful indicator to select the number of demonstrations to use during ICL.
Finally, we compare the ID curves from ICL and SFT directly and find that representations generated during ICL yield higher ID curves compared to fine-tuning on 1k samples, although SFT with 1k samples achieves the highest performance. This analysis provides evidence that the two learning paradims induce distinct representational structures in the embedding space, with representations created by ICL occupying higher-dimensional manifolds.
@@ 9 Limitations
In this study, we limit our analysis to models with sizes between 7B and 13B parameters. Future work may extend this line of investigation to models of larger sizes. We also focus on datasets defined by narrowly-defined tasks and do not consider datasets with long-form answers. Due to compute requirements, we only perform fine-tuning using LoRA adapters and do not study the impacts of full fine-tuning on intrinsic dimension. We also only analyze dense transformer models and do not study architectures such as Mixture-of-Experts (Shazeer et al., 2017) used in current frontier models.

### v2

@@ 7.1 Experiment Setup for Comparative Analysis
We conduct a series of experiments to directly compare the ID curves obtained from both SFT and ICL, following similar setups as discussed in Sections 5 and 6. For the fine-tuning experiments in this section, we train for only 4 epochs and measure the accuracy and ID solely at the final checkpoint. This choice is motivated by the observation in Section 5 that models tend to overfit beyond 4 epochs across the tested datasets.
For the ICL experiments, we consider values of $k\in\{0,1,2,5,10\}$ for the number of demonstrations. These values are popular in practice, and our previous experiments in Section 6 indicate that ID curves tend to plateau when $k\geq 10$. We perform these experiments on all 8 datasets and 4 models discussed in Section 4.
@@ 7.2 In-context Learning Induces Higher IDs Compared to Fine-Tuning
We find that across all datasets and models, ICL prompts with $k\geq 5$ consistently induces higher intrinsic dimensions (IDs) across all layers compared to both SFT and 0-shot prompts (see Figures 1b and 1c). This contrasts with the findings of Doimo et al. (2024), who find that SFT models often induces higher ID than models performing ICL.
@@
Figure 5: Boxplot displaying the distribution of normalized AUC values for different learning paradigms. Each point corresponds to the normalized AUC value for a (model, dataset) pair. The median normalized AUC peaks with 5-shot ICL, while values for SFT are closer to the 0-shot baseline (icl-0).
We also find that the ID values of models fine-tuned with 1000 samples tend to remain similar to the original ID of the baseline model on a zero-shot prompt (designated by icl-0). We present a heatmap displaying the average differences in normalized AUC between learning paradigms in Figure 4, and a boxplot depicting the distribution of normalized AUC values for the different paradigms in Figure 5.
@@ 7.3.1 Differing Shapes of Intrinsic Dimension Curves
We observe that the exact shape of the Intrinsic Dimension curves is highly dependent on the dataset. For some datasets, such as AG News, we observe a consistent "hunchback" shape, where the ID initially increases and then is progressively lower in the later layers of the model across all models and learning paradigms (Figure 36). This shape has been reported by previous work (Yin et al., 2024) in QA datasets. However, this pattern does not consistently hold across all models, datasets, and learning paradigms. For example, on the QQP dataset, we do not observe a consistent hunchback shape for icl-0, icl-1, or fine-tuning learning paradigms (Figure 33). In contrast, prior work has shown that Convolutional Neural Networks (Ansuini et al., 2019), as well as Image Generation Transformers such as ImageGPT and Protein Language Models (Valeriani et al., 2024), exhibit consistent Intrinsic Dimension patterns across their layers for inputs of their respective data modalities. This difference suggests that LLMs encode data into more diverse manifolds in their representation space, potentially reflecting their generality and the complexity of their learning tasks compared to other neural networks.
Figure 6: Boxplot displaying the distribution of normalized AUC values across datasets for each model in the ICL-5 shot setting. Each point corresponds to a (model, dataset) pair. The ID values lie in a narrow range, highlighting similarity in representation spaces across models.
We also find that, within a specific learning paradigm, the range of normalized AUC values across datasets is similar for the four different models we tested, despite the fact that these models come from different families and have different embedding dimensions (e.g., Llama-2-13b has a hidden dimension of 5120, while the other three models have hidden dimensions of 4096). Figure 6 depicts the range of normalized AUC values for the ICL-5 learning paradigm and shows that all values fall within a range of 20. We view this as evidence that different models may be generating representations with similar geometric complexity for a specific dataset, despite differences in model size or pre-training schemes. Similar boxplots for normalized AUC values from other experiments are included in Appendix E. These findings are in agreement with results from Cheng et al. (2023), which show that LLMs of different sizes and families create representations with similar ID values for a variety of text corpora.
@@ 7.4 Comparing Performance of Different Learning Paradigms
We found that models fine-tuned with 1k samples obtained the highest accuracy, while models performing ICL with 10 samples followed closely. This observation suggests that intrinsic dimension (ID) may not be directly related to accuracy: although fine-tuning with 1k samples yields ID values that remain closer to the baseline model, ICL models exhibit higher IDs yet achieve substantially lower accuracies. See Table 1 for the average performance of each learning paradigm across the models and datasets tested.
@@ 8 Summary
We present a detailed analysis of the intrinsic dimension (ID) induced by the SFT and ICL learning paradigms. Our experiments reveal that the normalized AUC of ID curves sometimes decreases during the initial stages of SFT but generally increases during the later stages.
Additionally, we observe that the normalized AUC of ID curves in ICL initially increases for small values of $k$ (the number of demonstrations) but plateaus or slightly decreases as $k$ increases further. Notably, the $k$ value corresponding to the highest normalized AUC also achieves peak (or near-peak) accuracy, suggesting that ID may serve as a useful indicator for selecting the optimal number of demonstrations during ICL.
Finally, our direct comparison of ID curves from ICL and SFT reveals that representations generated during ICL consistently yield higher ID curves compared to those from SFT on 1k samples, even though SFT with 1k samples achieves the highest overall performance. This analysis provides evidence that the two learning paradigms induce distinct representational structures in the embedding space, with ICL representations occupying higher-dimensional manifolds.
@@ 9 Limitations
In this study, we limit our analysis to models with sizes between 7B and 13B parameters. Future work may extend this investigation to models of different sizes. We also focus on datasets defined by narrowly focused tasks and do not consider datasets with long-form answers. Due to computational constraints, we perform fine-tuning only using LoRA adapters and do not explore the impacts of full fine-tuning on intrinsic dimension.